# Solutions of the singular Yamabe problem near singular boundaries

### v1

@@ Abstract.
In this paper, we investigate the asymptotic behaviors of solutions to the singular Yamabe problem with negative constant scalar curvature near singular boundaries and derive optimal estimates, where the background metrics are not assumed to be conformally flat. Specifically, we demonstrate that for a wide class of Lipschitz domains with asymptotic conical structure, the local positive solutions are well approximated by the positive solutions in the tangent cones at singular boundary points. This extends the results of [HJS, HS, SW].
@@ 1. Introduction
The singular Yamabe problem concerns finding a complete conformal metric of constant scalar curvature on a domain obtained by removing a nonempty closed subset from a compact Riemannian manifold. The singular Yamabe problem has been extensively investigated in a series of works including [AMC, LN, MaPa96, MaPa99, MPU96, MS91, SY88, Sc88]. In this paper, we focus on the asymptotic behaviors of solutions in the negative case, in which the resulting complete conformal metric has constant negative scalar curvature. For the asymptotic behaviors of solutions in the positive case, one may refer to [CGS1989, HLL2021, KMPS1999] (flat background metric) and [HXZ2023, M2008, XZ2022] (non-flat background metric).
Let $(M,g)$ be a smooth Riemannian manifold of dimension $n\geq 3$, which is either closed or non-compact and complete. Let $\Omega$ be a domain in $M$, and denote by $S_{g}$ the scalar curvature of $g$. Suppose $u>0$ solves
@@
then $u^{\frac{4}{n-2}}g$ is a complete metric with constant scalar curvature n(n1)-n(n-1) in $\Omega$. The existence theory for problem (1.1)–(1.2) was established in the works such as [AMC, LG, LN, SW]. In the case where the background manifold is the standard Euclidean space, (1.1)–(1.2) reduces to the Loewner-Nirenberg problem, namely,
@@
Let $V\subset\operatorname{\mathbb{R}}^{n}$ be an infinite cone with the origin being its vertex, and we assume $\Sigma=V\cap\mathbb{S}^{n-1}$ is a Lipschitz domain on $\mathbb{S}^{n-1}$. If $(M,g)=(\operatorname{\mathbb{R}}^{n},g_{0})$ and $\Omega=V\subset\operatorname{\mathbb{R}}^{n}$, where $g_{0}$ is the Euclidean metric, then there is a unique positive solution $u_{V}$ to (1.3)-(1.4). We note $u_{V}$ has the form
@@
When $\Omega$ is a sufficiently smooth domain, the asymptotic behaviors of positive solutions to (1.1)–(1.2) near the boundary has been thoroughly studied. For bounded $C^{2}$ domains $\Omega$, let $d$ denote the distance function to $\partial\Omega$ and $u$ be a positive solution of (1.3)-(1.4), Loewner and Nirenberg [LN] showed that near the boundary,
@@
where $C$ is a positive constant depending only on $n$ and $\Omega$. When $\partial\Omega$ is smooth, Mazzeo [M] and Andersson, Chruściel, and Friedrich [ACF] established an asymptotic expansion of solutions of (1.1)–(1.2) up to an arbitrarily finite order. The singular Yamabe problem has found numerous geometric applications in recent years, see [CMY2022, CLW2019, Graham2017, GG2021, GH2017] for example. In these works, the boundary behaviors of solutions to (1.1)–(1.2) play a crucial role.
Recently, the asymptotic behaviors of solutions (1.1)-(1.2) near singular boundaries have attracted growing interest. When the background metric is flat, Han and Shen [HS] proved that near a singular boundary point, the solution can be approximated by the corresponding solution in the tangent cone at that point. Besides, Jiang [Jiang] studied boundary expansion of solutions in finite cones over smooth spherical domains. Furthermore, Han, Jiang, and Shen [HJS] derived arbitrarily order asymptotic expansions of solutions in finite cones in terms of the corresponding solutions in infinite cones. For non-flat background metrics, Shen and Wang [SW] extended the results of [HS], where they applied a new method to construct barrier functions. In both [HS] and [SW], it is assumed that the boundary near a singular point consists of $k$ $(k\leq n)$ $C^{2}$ hypersurfaces intersecting at that point, with linearly independent normal vectors at that point. However, the asymptotic behaviors of solutions near more general types of singular points remain unknown.
@@ 1.1. Assumption
The following presents a type of singular boundary that differs from that in references [HS, SW].
@@
where $c$ is a small constant and $e_{n}=(0,\cdots,0,1)$. Set $T_{0}=S_{0}^{-1}$, $x=S_{0}y$ and $y=T_{0}x$ for $y\in V_{0}$ near 0. By a direct calculation,
@@
where $x^{\prime}=(x_{1},\cdots,x_{n-1},0)$. Then, $T_{0}(0)=0$, $\nabla T_{0}(0)=\operatorname{Id}$, and $\|T\|_{C^{2}}$ is bounded near 0. Let $\Omega_{0}=S_{0}(V_{0})$, then $\Omega_{0}$ is a domain with conical singularity at 0, and $V_{0}$ is the tangent cone of $\Omega_{0}$ at 0. Moreover, $\Omega_{0}$ can not be bounded by $k$ $C^{2}$-hypersurfaces near 0 with $k\leq n$.
@@ 1.2. Main results
Suppose $u>0$ in $\Omega\cap B_{1}(0)$ sovles
@@
Theorem 1.2 could be viewed as a generalization of the main results in [HS, SW]. Specifically, let $\partial\Omega\cap B_{1}(0)$ consist of $k$ $(k\leq n)$ $C^{2}$-hypersurfaces $S_{1},\cdots,S_{k}$ intersecting at 0 with the property that the normal vectors of $S_{1},\cdots,S_{k}$ at 0 are linearly independent. Then, the corresponding diffeomorphism $T$ can be explicitly constructed and satisfies (1.8)-(1.9), see [SW, Section 2]. Actually, estimate (1.12) is similar to (1.7) in some extent. When the background manifold is the standard Euclidean space and $\partial\Omega$ is $C^{2}$, the tangent cone at any boundary point is a half space, and for convenience we may assume the boundary point is 0 and the tangent cone is $\mathbb{R}^{n}_{+}$ in this case. For $x\in\Omega$ with $|x|$ small enough and $d(x,\partial\Omega)=d(x,0)$, estimate (1.12) at $x$ is actually the same as estimate (1.7) at $x$. From this perspective, the type of estimate in Theorem 1.2 is reasonable.
@@
Under the same assumptions as in Theorem 1.2, we further assume that $T\in C^{3}(B_{R}(0))$. Then, there is a constant $r_{1}>0$ small enough and a function $c_{1}\in C^{\infty}(\Sigma)\cap L^{\infty}(\Sigma)$ depending only on $n$, $g$, $V$, and the derivatives of $T$ at 0, such that
@@
for any $x\in\Omega\cap B_{r_{1}}(0)$, where
@@
with $\lambda_{1}$ represents the first eigenvalue of the elliptic operator $L_{1}$ defined by (A.8), and $C$ is a positive constant depending only on $n$, $g$, $\|T\|_{C^{3}(B_{R}(0))}$, and $V$.
@@
By Shen and Wang [SW, Proposition 4.5], when $n=3$ we know $\lambda_{1}>\frac{3}{4}$. Therefore, we have $\mu_{1}>1$.
@@
In fact, to obtain the term $c_{1}(Tx/|Tx|)|Tx|$ in (1.13), it suffices to assume that $T\in C^{2,\alpha}(B_{R}(0))$, for some positive constant $\alpha\in(0,1)$. Under this assumption, by an argument analogous to the proof of Theorem 1.3, we can obtain
$\left|\frac{u(x)}{u_{V}(Tx)}-1-c_{1}\left(\frac{Tx}{|Tx|}\right)|Tx|\right|\leq\begin{cases}\begin{aligned} &Cd^{1+\alpha}_{g}(x,0),\text{ if }\mu_{1}>1+\alpha,\\ &Cd^{1+\alpha}_{g}(x,0)|\log{d_{g}(x,0)}|,\text{ if }\mu_{1}=1+\alpha,\\ &Cd^{\mu_{1}}_{g}(x,0),\text{ if }\mu_{1}<1+\alpha.\end{aligned}\end{cases}$
Remark 1.6.
@@
Let $m$ be an integer with $1\leq m<\mu_{1}$. Under the same assumptions as in Theorem 1.2, we further assume that $T\in C^{m+1,\alpha}(B_{R}(0))$, for some positive constant $\alpha\in(0,1)$. Then, by a strategy similar to the proof of Theorem 1.3, we will find that there exist functions $c_{1},c_{2},\cdots,c_{m}\in C^{\infty}(\Sigma)\cap L^{\infty}(\Sigma)$ depending only on $n$, $g$, $V$, and the derivatives of $T$ at 0, such that
@@
$\left|\frac{u(x)}{u_{V}(Tx)}-1-\sum^{m}_{i=1}c_{i}\left(\frac{Tx}{|Tx|}\right)|Tx|^{i}\right|\leq\begin{cases}\begin{aligned} &Cd^{m+\alpha}_{g}(x,0),\text{ if }\mu_{1}>m+\alpha,\\ &Cd^{m+\alpha}_{g}(x,0)|\log{d_{g}(x,0)}|,\text{ if }\mu_{1}=m+\alpha,\\ &Cd^{\mu_{1}}_{g}(x,0),\text{ if }m<\mu_{1}<m+\alpha.\end{aligned}\end{cases}$
Note that when $T=\operatorname{Id}$ and the background manifold is the standard Euclidean space with $\Omega=V\cap B_{2}$, Han, Jiang, and Shen [HJS] obtained arbitrarily high-order asymptotic expansions for $u(x)/u_{V}(x)$. However, their discussion relies heavily on the fact that both $u$ and $u_{V}$ satisfy the same equation. Specifically, they consider the equation satisfied by $|x|^{(n-2)/2}(u-u_{V})(x)$. In this scenario, the problem can be elegantly transformed into the asymptotic analysis of an ordinary differential equation with respect to the radial variable, coupled with establishing the Fredholm theory for a certain singular elliptic operator on Lipschitz domains of the sphere. When $T$ is a small perturbation of the identity map, the equation governing $|y|^{(n-2)/2}(u\circ T^{-1}-u_{V})(y)$ will generate perturbation terms involving $u_{i},u_{ij}$. Consequently, due to the lack of precise estimates for $u_{i},u_{ij}$, their method is no longer applicable in this case. On the other hand, the coefficient functions on the sphere in the expansions of [HJS] satisfy a decay of $O(d^{\tau}_{\Sigma})$, which aligns with the Fredholm theory of the singular elliptic operator $L_{1}$ defined by (A.8). Here, $d_{\Sigma}$ is the distance function to $\partial\Sigma$ on the sphere and $\tau$ is a positive constant. However, in our more general setting, we can only guarantee that $c_{i}$ are bounded functions and cannot expect them to exhibit decay near the boundary, as demonstrated by Example 5.2 and Example 5.3.
@@
This paper is organized as follows. Section 2 presents some preliminary results. Section 3 is devoted to the proof of Theorem 1.2. In Section 4, we prove an existence result for uniformly degenerate elliptic equations in Lipschitz domains, which is then applied to prove Theorem 1.3. We construct some examples with no tangential decay in Section 5, which guarantee that estimate (1.12) is optimal. Finally, in the appendix (Section A), we discuss some existence results and estimates for degenerate or singular elliptic equations that are crucial to the proof of Theorem 1.3.
@@
Acknowledgements: Weiming Shen is supported by NSFC Grant 12371208. Zhehui Wang is supported in part by NSFC Grant 12401247, and Guangdong Basic and Applied Basic Research Foundation (Grant No. 2023A1515110910). Jiongduo Xie is partially supported by NSFC Grant W2531006, NSFC Grant 12250710674, NSFC Grant 12031012, and the Center of Modern Analysis-A Frontier Research Center of Shanghai.
@@ 2. Preliminaries
This section provides some preliminary results needed for our analysis. The following error estimates for the derivatives of a certain function under a $C^{2}$-diffeomorphism $T$ will be applied in the proof in the case when $n=3$.
@@
Let $\Omega$, $V$, and $T$ be as in assumption 1.1. Then, for any $f\in C^{2}(V\cap T(B_{R}(0)))$ and $x\in\Omega\cap B_{R}(0)$, we have,
@@
where $C_{0}$ is a constant depending only on $C^{2}$-norm of $T$.
@@
Let $\varphi(x)=Tx-x$ and set $\varphi(x)=(\varphi^{1}(x),\cdots,\varphi^{n}(x))$. By assumptions we have
@@
we have,
$|\partial_{x_{i}}(f\circ T(x))-\partial_{y_{i}}f(Tx)|$ $=|\partial_{y_{k}}f(Tx)\partial_{x_{i}}\varphi^{k}(x)|$
$\leq|\nabla f(Tx)|\|\nabla^{2}T\|_{\infty}|x|.$
Further, by (2.3), $\nabla^{2}\varphi=\nabla^{2}T$, and
@@
$\qquad\qquad\qquad\qquad\qquad+\partial_{y_{k}}f(Tx)\partial_{x_{i}x_{j}}^{2}\varphi^{k}(x),$
we have,
@@
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad+|\partial_{y_{k}}f(Tx)\partial_{x_{i}x_{j}}^{2}\varphi^{k}(x)|$
$\leq$ $|\nabla^{2}f(Tx)|\|\nabla^{2}T\|_{\infty}|x|+|\nabla^{2}f(Tx)|(1+\|\nabla^{2}T\|_{\infty}|x|)\|\nabla^{2}T\|_{\infty}|x||$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad+|\nabla f(Tx)|\|\nabla^{2}T\|_{\infty}$
@@
Let $\xi\in C^{\infty}(\Sigma)$ be a positive function in $\Sigma$ satisfying (1.5)-(1.6). By [HJS, Lemma 2.4 and Lemma 2.5], the functions $\xi$ and $\rho=\xi^{-\frac{2}{n-2}}$ have the following property.
@@
Let $\Sigma\subset\mathbb{S}^{n-1}$ be a Lipschitz domain, $d_{\Sigma}=d_{g_{\mathbb{S}^{n-1}}}(\cdot,\partial\Sigma)$, and $\xi\in C^{\infty}(\Sigma)$ be a positive function in $\Sigma$ satisfying (1.5)-(1.6). Then,
@@
For $x\in B_{R}(0)$, set $y=Tx$, and $v(y)=u(T^{-1}y)$. By a direct calculation, $v$ satisfies the following equation:
$\mathcal{L}v:=a_{ij}v_{ij}+b_{i}v_{i}+cv=\frac{n(n-2)}{4}v^{\frac{n+2}{n-2}},\quad\mbox{ in }V\cap T(B_{R}(0)),$
@@
and $T\in C^{2}$ with $\nabla T(0)=\operatorname{Id}$, then, $a_{ij}\in C^{1}$, $b_{i}\in C^{0}$ and
@@
and $r=|y|$. For positive constants $A$ and $B$ to be determined, we set
@@
$D_{1}$ $=-\frac{n(n+2)}{4}(1-\frac{n-2}{n+2}\beta)Au_{V}^{\beta+\frac{4}{n-2}}+A\beta(\beta-1)u_{V}^{\beta-2}|\nabla u_{V}|^{2},$
$D_{2}$ $=-\frac{n(n+2)}{4}(1-\frac{n-2}{n+2})Bru_{V}^{\frac{n+2}{n-2}},$
$D_{3}$ $=2B\nabla r\cdot\nabla u_{V}+B\frac{n-1}{r}u_{V},$
and by (3.1) we note
@@
According to [HS, Remark 3.3, Lemma 3.4], we know there is constant $C_{1}>1$ depending only on $n$ and the size of exterior cones such that for $y\in V\cap T(B_{R}(0))$,
(3.2) $d^{2}(y)|\nabla^{2}u_{V}(y)|+d(y)|\nabla u_{V}(y)|\leq C_{1}u_{V}(y)$
@@
(3.4) $|D_{4}|\leq C_{2}\frac{u_{V}r}{d^{2}}(1+A\beta u_{V}^{\beta-1}+Br),$
where $C_{2}>0$ is a constant depending only on $n$ and the size of exterior cones. By $0\leq\beta<1$, we have
$D_{1}\leq-\frac{n(n+2)}{4}(1-\frac{n-2}{n+2}\beta)Au_{V}^{\beta+\frac{4}{n-2}}=\left\{\begin{aligned} -\frac{n(n+2)}{4}(1-\frac{n-2}{n+2}\beta)Au_{V}u_{V}^{\frac{2}{n-2}},&\quad n\geq 4,\\ -\frac{n(n+2)}{4}(1-\frac{n-2}{n+2}\beta)Au_{V}u_{V}^{3},&\quad n=3.\end{aligned}\right$.
@@
(3.5) $D_{1}\leq-\frac{n(n+2)}{4}(1-\frac{n-2}{n+2}\beta)\frac{C_{1}^{-\frac{2}{n-2}}Au_{V}}{d},\quad\mbox{ for }y\in V\cap T(B_{R}(0))\cap B_{r_{1}}(0).$

### v2

@@ Abstract.
In this paper, we investigate the asymptotic behaviors of solutions to the singular Yamabe problem with negative constant scalar curvature near singular boundaries and derive optimal estimates, where the background metrics are not assumed to be conformally flat. Specifically, we demonstrate that for a wide class of Lipschitz domains with asymptotic conical structure, the local positive solutions are well approximated by the positive solutions in the tangent cones at singular boundary points. This extends the results of [10, 12, 20].
@@ 1. Introduction
The singular Yamabe problem concerns finding a complete conformal metric of constant scalar curvature on a domain obtained by removing a nonempty closed subset from a compact Riemannian manifold. In this paper, we focus on the asymptotic behaviors of solutions in the case that the resulting complete conformal metric has constant negative scalar curvature.
Let $(M,g)$ be a smooth Riemannian manifold of dimension $n\geq 3$, which is either closed or non-compact and complete. Let $\Omega\subset M$ be a domain, and denote by $S_{g}$ the scalar curvature of $g$. Suppose $u>0$ solves
@@
then $u^{\frac{4}{n-2}}g$ is a complete metric with constant scalar curvature n(n1)-n(n-1) in $\Omega$. The existence theory for problem (1.1)–(1.2) was established in the works such as [2, 16, 17, 20]. In the case where the background manifold is the standard Euclidean space, (1.1)–(1.2) is reduced to the Loewner-Nirenberg problem, namely,
@@
Let $V\subset\operatorname{\mathbb{R}}^{n}$ be an infinite cone with the origin being its vertex, and we assume $\Sigma=V\cap\mathbb{S}^{n-1}$ is a Lipschitz domain on $\mathbb{S}^{n-1}$. If $(M,g)=(\operatorname{\mathbb{R}}^{n},g_{0})$ and $\Omega=V\subset\operatorname{\mathbb{R}}^{n}$, where $g_{0}$ is the Euclidean metric, then there is a unique positive solution $u_{V}$ to (1.3)-(1.4). From [10, Theorem 2.3], $u_{V}$ has the form
@@
When $\Omega$ is a sufficiently smooth domain, the asymptotic behaviors of positive solutions to (1.1)–(1.2) near the boundary has been thoroughly studied. For bounded $C^{2}$ domains $\Omega$, let $d$ denote the distance function to $\partial\Omega$ and let $u$ be a positive solution of (1.3)-(1.4), Loewner and Nirenberg [17] showed that near the boundary,
@@
where $C$ is a positive constant depending only on $n$ and $\Omega$. When $\partial\Omega$ is smooth, Mazzeo [19] and Andersson, Chruściel, and Friedrich [1] established asymptotic expansions of solutions of (1.1)–(1.2) up to an arbitrarily finite order. The singular Yamabe problem has found numerous geometric applications in recent years, see [4, 5, 8, 6, 9] for example. In these works, the boundary behaviors of solutions to (1.1)–(1.2) play a crucial role.
Recently, the asymptotic behaviors of solutions (1.1)-(1.2) near singular boundaries have attracted increasing interest. When the background metric is flat, Han and Shen [12] proved that near a singular boundary point, the solution can be approximated by the corresponding solution in the tangent cone at that point. Besides, Jiang [14] studied boundary expansion of solutions in finite cones over smooth spherical domains. Furthermore, Han, Jiang, and Shen [10] derived arbitrarily finite order asymptotic expansions of solutions in finite cones in terms of the corresponding solutions in infinite cones. For non-flat background metrics, Shen and Wang [20] extended the results of [12], where they applied a new method to construct barrier functions. In both [12] and [20], it is assumed that the boundary near a singular point consists of $k$ $(k\leq n)$ $C^{2}$-hypersurfaces intersecting at that point, with linearly independent normal vectors at that point. However, the asymptotic behaviors of solutions near more general types of singular points remain unknown. Regarding the asymptotic behaviors of solutions of the singular Yamabe problem with positive constant scalar curvature, relevant results can be found in [3, 11, 15] (flat background metric) and in [13, 18, 21] (non-flat background metric).
@@ 1.1. Assumption
The following presents a type of singular boundary that differs from that in references [12, 20].
@@
where $c>0$ is a small constant and $e_{n}=(0,\cdots,0,1)$. Set $T_{0}=S_{0}^{-1}$, $x=S_{0}y$, and $y=T_{0}x$ for $y\in V_{0}$ near 0. By a direct calculation,
@@
where $x^{\prime}=(x_{1},\cdots,x_{n-1},0)$. Then, $T_{0}(0)=0$, $\nabla T_{0}(0)=\operatorname{Id}$, and $\|T\|_{C^{2}}$ is bounded near 0. Let $\Omega_{0}=S_{0}(V_{0})$, then $\Omega_{0}$ is a domain with conical singularity at 0, and $V_{0}$ is the tangent cone of $\Omega_{0}$ at 0. Moreover, near 0, $\partial\Omega_{0}$ does not consist of $k$ $C^{2}$-hypersurfaces intersecting at 0 for any $k\leq n$.
@@ 1.2. Main results
Suppose $u>0$ in $\Omega\cap B_{1}(0)$ solves
@@
Theorem 1.2 could be viewed as a generalization of the main results in [12, 20]. Specifically, let $\partial\Omega\cap B_{1}(0)$ consist of $k$ $(k\leq n)$ $C^{2}$-hypersurfaces $S_{1},\cdots,S_{k}$ intersecting at 0 with the property that the normal vectors of $S_{1},\cdots,S_{k}$ at 0 are linearly independent. Then, the corresponding diffeomorphism $T$ can be explicitly constructed and satisfies (1.8)-(1.9), see [20, Section 2]. Actually, estimate (1.12) is similar to (1.7) in some extent. When the background manifold is the standard Euclidean space and $\partial\Omega$ is $C^{2}$, the tangent cone at any boundary point is a half space, and for convenience we may assume the boundary point is 0 and the tangent cone is $\mathbb{R}^{n}_{+}$ in this case. For $x\in\Omega$ with $|x|$ small enough and $d(x,\partial\Omega)=d(x,0)$, estimate (1.12) at $x$ is similar to estimate (1.7) at $x$. From this perspective, the type of estimate in Theorem 1.2 is reasonable.
@@
Under the same assumptions as in Theorem 1.2, we further assume that $T\in C^{3}(B_{R}(0))$. Then, there is a constant $r_{1}>0$ small enough and a function $c_{1}\in C^{\infty}(\Sigma)\cap L^{\infty}(\Sigma)$, such that
@@
for any $x\in\Omega\cap B_{r_{1}}(0)$, where $c_{1}$ depends only on $n$, $g$, $V$, and the derivatives of $T$ at 0,
@@
with $\lambda_{1}$ represents the first eigenvalue of the elliptic operator $L_{1}$ defined by (A.8) depending only on $n$ and $\Sigma$, and $C$ is a positive constant depending only on $n$, $g$, $\|T\|_{C^{3}(B_{R}(0))}$, and $V$. In addition, when $\mu_{1}=2$ and $T\in C^{3,\gamma}(B_{R}(0))$ for some $\gamma\in(0,1)$, we can improve the estimate as follows:
(1.15) $\left|\frac{u(x)}{u_{V}(Tx)}-1-c_{1}\left(\frac{Tx}{|Tx|}\right)|Tx|-c_{2}\left(\frac{Tx}{|Tx|}\right)|Tx|^{2}\log|Tx|\right|\leq Cd^{2}_{g}(x,0),$
where $c_{2}\in C^{\infty}(\Sigma)\cap L^{\infty}(\Sigma)$ depends only on $n$, $g$, $V$, and the derivatives of $T$ at 0.
@@
By Shen and Wang [20, Proposition 4.5], when $n=3$ we know $\lambda_{1}>\frac{3}{4}$. Therefore, we have $\mu_{1}>1$.
It should be mentioned here that, by formal calculations, we can not determine the coefficient of the term with order $r^{\mu_{1}}$. See Remark 4.2 for a detailed discussion.
@@
In fact, to obtain the term $c_{1}(Tx/|Tx|)|Tx|$ in (1.13), it suffices to assume that $T\in C^{2,\alpha}(B_{R}(0))$. In general, let $m$ be an integer with $1\leq m\leq\mu_{1}$. Under the same assumptions as in Theorem 1.2, we further assume that $T\in C^{m+1,\alpha}(B_{R}(0))$, for some positive constant $\alpha\in[0,1]$ with $m+\alpha>1$. Here, $C^{m,0}$ denotes $C^{m}$. Then, by a strategy similar to the proof of Theorem 1.3, we will find that there exist functions $c_{1},c_{2},\cdots,c_{m}\in C^{\infty}(\Sigma)\cap L^{\infty}(\Sigma)$ depending only on $n$, $g$, $V$, $m$, $\alpha$, and the derivatives of $T$ at 0, such that
@@
$\left|\frac{u(x)}{u_{V}(Tx)}-1-\sum^{m-[1-\alpha]}_{i=1}c_{i}\left(\frac{Tx}{|Tx|}\right)|Tx|^{i}\right|\leq\begin{cases}\begin{aligned} &Cd^{m+\alpha}_{g}(x,0),\text{ if }\mu_{1}>m+\alpha,\\ &Cd^{m+\alpha}_{g}(x,0)|\log{d_{g}(x,0)}|,\text{ if }\mu_{1}=m+\alpha,\\ &Cd^{\mu_{1}}_{g}(x,0),\text{ if }m<\mu_{1}<m+\alpha,\end{aligned}\end{cases}$
and if $\mu_{1}=m$ and $\alpha>0$, we have
$\left|\frac{u(x)}{u_{V}(Tx)}-1-\sum^{m-1}_{i=1}c_{i}\left(\frac{Tx}{|Tx|}\right)|Tx|^{i}-c_{m}\left(\frac{Tx}{|Tx|}\right)|Tx|^{m}\log|Tx|\right|\leq Cd^{m}_{g}(x,0).$
Note that when $T=\operatorname{Id}$ and the background manifold is the standard Euclidean space with $\Omega=V\cap B_{2}$, Han, Jiang, and Shen [10] proved
(1.16) $\left|\frac{u(x)}{u_{V}(x)}-1\right|\leq Cd_{\Sigma}^{\tau}|x|^{\mu_{1}},$
for some constant $\tau>0$, where $d_{\Sigma}$ is the distance function to $\partial\Sigma$ on the sphere. They also obtained arbitrarily high-order asymptotic expansions for $u(x)/u_{V}(x)$. There are two key differences between our estimates and (1.16):
The first difference is that the expansion of $u/u_{V}$ in (1.16) does not have terms with order greater than 0 and less than $\mu_{1}$. We point out that the discussion in [10] relies heavily on the fact that both $u$ and $u_{V}$ satisfy the same equation. The second difference is that the coefficient functions on $\Sigma$} in (1.16) and the high-order expansions of [10] satisfy a decay of $O(d^{\tau}_{\Sigma})$, which aligns with the Fredholm theory in the $L^{2}$-setting of the singular elliptic operator $L_{1}$ defined by (A.8), and it is essential for them to derive arbitrarily high-order asymptotic expansions. However, in our more general setting, $T$ is a perturbation of the identity map, and we can only guarantee that $c_{i}$ are bounded functions and cannot expect them to exhibit decay near the boundary, as demonstrated by Example 5.2 and Example 5.3. Example 5.3 also implies that the term with order $1$ does indeed appear. Using the inner-outer gluing method, we prove a general existence result (Lemma 4.1) for a class of uniformly degenerate elliptic equations in bounded Lipschitz domains. A key novelty of our existence result is that it allows the zero-order coefficient to be positive, or the non-homogeneous term to be non-$L^{2}$.
@@
This paper is organized as follows. Section 2 presents some preliminary results. Section 3 is devoted to the proof of Theorem 1.2. In Section 4, we prove an existence result for uniformly degenerate elliptic equations in Lipschitz domains, which is then applied to prove Theorem 1.3. We construct some examples with no tangential decay in the left-hand side of (1.12) in Section 5, which guarantee that estimate (1.12) is optimal. Finally, in the appendix (Section A), we discuss some existence results and estimates for degenerate or singular elliptic equations that are crucial to the proof of Theorem 1.3.
@@
Acknowledgements: Weiming Shen is supported by NSFC Grant 12371208. Zhehui Wang is supported in part by NSFC Grant 12401247 and Guangdong Basic and Applied Basic Research Foundation (Grant No. 2023A1515110910). Jiongduo Xie is partially supported by NSFC Grant W2531006, NSFC Grant 12250710674, NSFC Grant 12031012, and the Center of Modern Analysis-A Frontier Research Center of Shanghai.
@@ 2. Preliminaries
This section provides some preliminary results needed for the subsequent analysis. The following error estimates for the derivatives of a certain function under a $C^{2}$-diffeomorphism $T$ will be applied in the proof in the case $n=3$.
@@
Let $\Omega$, $V$, and $T$ be as in assumption 1.1. Then, for any $f\in C^{2}(V\cap T(B_{R}(0)))$ and $x\in\Omega\cap B_{R}(0)$, we have
@@
where $C_{0}$ is a constant depending only on the $C^{2}$-norm of $T$.
@@
Let $\varphi(x)=Tx-x$ and set $\varphi(x)=(\varphi^{1}(x),\cdots,\varphi^{n}(x))$. By assumptions, we have
@@
we have
$|\partial_{x_{i}}(f\circ T(x))-\partial_{y_{i}}f(Tx)|$ $=|\partial_{y_{k}}f(Tx)\partial_{x_{i}}\varphi^{k}(x)|\leq|\nabla f(Tx)|\|\nabla^{2}T\|_{\infty}|x|.$
Furthermore, by (2.3), $\nabla^{2}\varphi=\nabla^{2}T$, and
@@
$\qquad+\partial_{y_{k}}f(Tx)\partial_{x_{i}x_{j}}^{2}\varphi^{k}(x),$
we have
@@
$\qquad+|\partial_{y_{k}}f(Tx)\partial_{x_{i}x_{j}}^{2}\varphi^{k}(x)|$
$\leq|\nabla^{2}f(Tx)|\|\nabla^{2}T\|_{\infty}|x|+|\nabla^{2}f(Tx)|(1+\|\nabla^{2}T\|_{\infty}|x|)\|\nabla^{2}T\|_{\infty}|x|$
$\qquad+|\nabla f(Tx)|\|\nabla^{2}T\|_{\infty}$
@@
Let $\xi\in C^{\infty}(\Sigma)$ be a positive function in $\Sigma$ satisfying (1.5)-(1.6). By [10, Lemma 2.4 and Lemma 2.5], the functions $\xi$ and $\rho=\xi^{-\frac{2}{n-2}}$ have the following property.
@@
Let $\Sigma\subsetneq\mathbb{S}^{n-1}$ be a Lipschitz domain, $d_{\Sigma}=d_{g_{\mathbb{S}^{n-1}}}(\cdot,\partial\Sigma)$, and $\xi\in C^{\infty}(\Sigma)$ be a positive function in $\Sigma$ satisfying (1.5)-(1.6). Then,
@@
For $x\in B_{R}(0)$, set $y=Tx$ and $v(y)=u(T^{-1}y)$. By a direct calculation, $v$ satisfies the following equation:
$\mathcal{L}v:=a_{ij}v_{ij}+b_{i}v_{i}+cv=\frac{n(n-2)}{4}v^{\frac{n+2}{n-2}}\quad\mbox{ in }V\cap T(B_{R}(0)),$
@@
and $T\in C^{2}$ with $\nabla T(0)=\operatorname{Id}$, then $a_{ij}\in C^{1}$, $b_{i}\in C^{0}$ and
@@
and $r=|y|$. For some positive constants $A$ and $B$ to be determined, we set
@@
$D_{1}$ $=-\frac{n(n+2)}{4}\left(1-\frac{n-2}{n+2}\beta\right)Au_{V}^{\beta+\frac{4}{n-2}}+A\beta(\beta-1)u_{V}^{\beta-2}|\nabla u_{V}|^{2},$
$D_{2}$ $=-\frac{n(n+2)}{4}\left(1-\frac{n-2}{n+2}\right)Bru_{V}^{\frac{n+2}{n-2}},$
$D_{3}$ $=2B\nabla r\cdot\nabla u_{V}+B\frac{n-1}{r}u_{V}.$
By (3.1), we know
@@
According to [12, Remark 3.3, Lemma 3.4], we know that there is a constant $C_{1}>1$ depending only on $n$ and the size of the exterior cones such that for $y\in V\cap T(B_{R}(0))$,
(3.2) $d^{2}(y)|\nabla^{2}u_{V}(y)|+d(y)|\nabla u_{V}(y)|\leq C_{1}u_{V}(y),$
@@
(3.4) $|D_{4}|\leq C_{2}\frac{u_{V}r}{d^{2}}(1+A(\beta+1)u_{V}^{\beta-1}+Br),$
where $C_{2}>0$ is a constant depending only on $n$ and the size of the exterior cones. Note $0\leq\beta<1$, we have
$D_{1}\leq-\frac{n(n+2)}{4}\left(1-\frac{n-2}{n+2}\beta\right)Au_{V}^{\beta+\frac{4}{n-2}}=\left\{\begin{aligned} -\frac{n(n+2)}{4}\left(1-\frac{n-2}{n+2}\beta\right)Au_{V}u_{V}^{\frac{2}{n-2}},&\quad n\geq 4,\\ -\frac{n(n+2)}{4}\left(1-\frac{n-2}{n+2}\beta\right)Au_{V}u_{V}^{3},&\quad n=3.\end{aligned}\right$.
@@
(3.5) $D_{1}\leq-\frac{n(n+2)}{4}\left(1-\frac{n-2}{n+2}\beta\right)\frac{C_{1}^{-\frac{2}{n-2}}Au_{V}}{d},\quad\mbox{for }y\in V\cap T(B_{R}(0))\cap B_{r_{1}}(0).$